# Quantum cohomology of projective bundles

### v3

@@ 1. Introduction
For a rank-$r$ vector bundle $V\to B$ over a smooth projective variety $B$, we can construct the projective bundle $\mathbb{P}(V)\to B$ consisting of lines in the fibres of $V\to B$. The ordinary cohomology group $H^{*}(\mathbb{P}(V))$ of $\mathbb{P}(V)$ is generated by the relative hyperplane class $p:=c_{1}(\mathcal{O}_{\mathbb{P}(V)}(1))$ as an $H^{*}(B)$-algebra as follows:
@@
where $q$ is the Novikov variable associated with vertical curves. It decomposes into a direct sum of $H^{*}(B)$ as a ring for $q\neq 0$. This decomposition must persist under full quantum deformation, parametrized by $q$, the Novikov variable $Q$ of $B$ and the bulk parameter $\hat{\tau}\in H^{*}(\mathbb{P}(V))$. The question then arises as to whether each summand is isomorphic to the quantum cohomology of $B$. This paper aims to affirm this correspondence, i.e. $QH^{*}(\mathbb{P}(V))$ decomposes into a direct sum of $QH^{*}(B)$. Furthermore, we show that the decomposition can be lifted to quantum cohomology $D$-modules.
@@ 1.1. Mirror theorem for non-split projective bundles
The equivariant $J$-function $J_{V}^{\lambda}$ is defined over the Novikov ring of $B$. We use the splitting of $H_{2}(\mathbb{P}(V),\mathbb{Z})\twoheadrightarrow H_{2}(B,\mathbb{Z})$ given by the class $p=c_{1}(\mathcal{O}(1))$ to identify Novikov variables for $B$ with Novikov variables for $\mathbb{P}(V)$, see Notation 3.1.
@@
The quantum Riemann-Roch theorem of Coates-Givental [7] determines the $S^{1}$-equivariant $J$-function $J_{V}^{\lambda}$ from the genus-zero Gromov-Witten invariants of the base $B$ and the Chern classes $c_{i}(V)$. Therefore Theorem 1.1 determines all genus-zero Gromov-Witten invariants of $\mathbb{P}(V)$ from those of the base $B$. This recovers the genus-zero part of the results of Fan-Lee [16] and Fan [15], who showed that all genus Gromov-Witten invariants of $\mathbb{P}(V)$ can be reconstructed from those of the base $B$ and the Chern classes $c_{i}(V)$. On the other hand, in contrast to the split case, it is not easy to find a closed formula for $J_{V}^{\lambda}$ in general. We give a non-split example, the cotangent bundle of $\mathbb{P}^{n}$, in Example 3.5.
@@
For other techniques to compute Gromov-Witten invariants of non-split projective bundles, we refer the reader to the reconstruction algorithm [49] and the quantum splitting principle [36, 14].
@@ 1.2. Decomposition of quantum $D$-modules
An important property of quantum cohomology is that it forms a $D$-module over the parameter space. Roughly speaking, the quantum $D$-module $\operatorname{QDM}(X)$ of a smooth projective variety $X$ is the module $H^{*}(X)[z][\![Q,\tau]\!]$ equipped with the flat covariant derivatives $\nabla_{\tau^{i}}$, $\nabla_{z\partial_{z}}$, $\nabla_{\xi Q\partial_{Q}}$ with respect to the parameters $\tau$, $z$, $Q$ (see §2.2). Here $\tau\in H^{*}(X)$ represents the bulk parameter and $Q$ denotes the Novikov variable. These covariant derivatives, called the quantum connection, are defined in terms of the quantum multiplication. The quantum $D$-module is also equipped with a $z$-sesquilinear pairing $P$, induced from the Poincaré pairing and compatible with the quantum connection. The $I$- or $J$-functions appearing in the previous section are cohomology-valued solutions of the quantum $D$-module.
@@
The Mirror Theorem 1.1 establishes that the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ is a Fourier transform of the $S^{1}$-equivariant quantum $D$-module $\operatorname{QDM}_{S^{1}}(V)$ (see Theorem 1.11 below). Using this fact together with the quantum Riemann-Roch (QRR) theorem of Coates-Givental [7], which connects the quantum $D$-modules of $V$ and $B$, we derive the following decomposition theorem for $\operatorname{QDM}(\mathbb{P}(V))$.
$Q$$q^{-c_{1}(V)/r}Q$$q$$q^{-1/r^{\prime}}$(($\operatorname{QDM}(\mathbb{P}(V))$$\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)$
@@
Figure 1. Decomposition pictured over the Kähler moduli space: $\operatorname{QDM}(\mathbb{P}(V))$ splits into a direct sum of $\operatorname{QDM}(B)$ near the infinity divisor $(q^{-1/r^{\prime}}=0)$, where $Q$ is the Novikov variable of the base $B$ and $q$ is the Novikov variable of the fibre $\mathbb{P}^{r-1}$.
Theorem 1.7 (Theorem 5.1; see also Figure 1).
@@
Let $V\to B$ be as in Theorem 1.1. We have a map $\varsigma=\bigoplus_{j=0}^{r-1}\varsigma_{j}\colon H^{*}(\mathbb{P}(V))\to\bigoplus_{j=0}^{r-1}H^{*}(B)$ and a decomposition of the quantum $D$-modules
$\Phi\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{i=0}^{r-1}\varsigma_{i}^{*}\operatorname{QDM}(B)_{\rm loc}$
where $\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}$ and $\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ denote the localized quantum $D$-modules of $\mathbb{P}(V)$ and $B$ defined over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$, where $Q$ is the Novikov variable of the base $B$ and $r^{\prime}$ is in (5.1).
Corollary 1.8.
The derivative of the map $\varsigma$ induces an isomorphism of the quantum cohomology rings $(H^{*}(\mathbb{P}(V)),\star_{\hat{\tau}})\cong\bigoplus_{i=0}^{r-1}(H^{*}(B),\star_{\varsigma_{i}(\hat{\tau})})$. In particular, the quantum cohomology of $\mathbb{P}(V)$ is generically semisimple if and only if the same holds for the base $B$.
@@
In the split case, a similar decomposition of the quantum $D$-module (more generally, for toric bundles) follows from Brown’s theorem [3]: this has been discussed by Coates-Givental-Tseng [8] and Koto [34]. These previous works rely on the equivariant localization with respect to the (×$(\mathbb{C}^{\times})^{r-1}$-action (which exists only in the split case) and our argument is substantively different from them.
@@ 1.3. Equivariant quantum cohomology and Fourier transformation
$[\lambda,\mathbb{S}]=z\mathbb{S}.$
@@
We can rephrase the mirror theorem (Theorems 1.1, 3.3) as the Fourier duality between the quantum $D$-modules of $V$ and $\mathbb{P}(V)$, as follows:
@@ Acknowledgements
H.I. thanks Fumihiko Sanda, Yuuji Tanaka and Constantin Teleman for very useful discussions on shift operators and equivariant cohomology. H.I. is supported by JSPS grant 16H06335, 20K03582, 21H04994 and 23H01073. Y.K. is supported by JSPS grant 21J23023.
@@ 2.1. Gromov-Witten invariants and quantum cohomology
A reference on Gromov-Witten invariants is made to [40]. Let $X$ be a smooth projective variety over $\mathbb{C}$. For $n\geq 0$ and $d\in H_{2}(X,\mathbb{Z})$, let $X_{0,n,d}$ denote the moduli stack of genus-zero, $n$-points, degree $d$ stable maps to $X$. This is a proper Deligne-Mumford stack carrying a virtual fundamental class $[X_{0,n,d}]_{\rm vir}\in H_{2D}(X_{0,n,d},\mathbb{Q})$ of degree $D=\dim_{\mathbb{C}}X+c_{1}(TX)\cdot d+n-3$. Let $\operatorname{ev}_{i}\colon X_{0,n,d}\to X$ be the evaluation map at the $i$th marked point. The (genus-zero, descendant) Gromov-Witten invariants are the integrals of the form:
@@
Let $\operatorname{Eff}(X)\subset H_{2}(X,\mathbb{Z})$ denote the semigroup of effective curve classes. For a module $K$, we write $K[\![Q]\!]$ for the space of formal power series $\sum_{d\in\operatorname{Eff}(X)}k_{d}Q^{d}$ with $k_{d}\in K$. The formal variable $Q$ is called the Novikov variable. The module $K[\![Q]\!]$ has a natural ring structure when $K$ is a ring. The quantum cohomology will be defined over the ring $\mathbb{C}[\![Q]\!]$, which is called the Novikov ring.
@@
for $\alpha,\beta,\gamma\in H^{*}(X)$. The product $\star_{\tau}$ introduces an associative and supercommutative ring structure on $H^{*}(X)[\![Q,\tau]\!]=H^{*}(X)[\![Q]\!][\![\tau^{0},\dots,\tau^{s}]\!]$ such that $1\in H^{0}(X)$ is the unity. This is called quantum cohomology. The variables $\tau^{i}$ associated with odd classes $\phi_{i}$ have odd parity, and therefore anti-commute with each other by supercommutativity; more precisely we have $\tau^{i}\tau^{j}=(-1)^{|i||j|}\tau^{j}\tau^{i}$, $\tau^{i}\phi_{j}=(-1)^{|i||j|}\phi_{j}\tau^{i}$ with $|i|=\deg\phi_{i}$. The Novikov variable $Q$ is even and commutes with other elements. We also follow the Koszul sign convention when expanding the correlator $\left\langle\alpha,\beta,\gamma,\tau,\dots,\tau\right\rangle_{0,n+3,d}^{X}$ in power series of $\tau^{i}$. The parameter space $H^{*}(X)$ should be viewed as a formal supermanifold over $\mathbb{C}[\![Q]\!]$; we refer to [40] for the details.
@@ 2.2. Quantum connection and fundamental solution
The quantum connection admits the fundamental solution $M_{X}(\tau)$ defined as follows (see [19, §1], [44, Proposition 2], [25, Proposition 2.4], [6, Proposition 3.1] for the properties below):
@@
where $\{\phi^{j}\}$ is the dual basis of $\{\phi_{i}\}$ such that $(\phi_{i},\phi^{j})_{X}=\delta_{i}^{j}$. We used the String Equation [44, §1.2] in the second line. The other columns $M_{X}(\tau)\phi_{i}$ of $M_{X}(\tau)$ are given by the derivatives $z\partial_{\tau^{i}}J_{X}(\tau)$ of the $J$-function due to the differential equation (2.7).

### v4

@@ 1. Introduction
For a rank-$r$ vector bundle $V\to B$ over a smooth projective variety $B$, we can construct the projective bundle $\pi\colon\mathbb{P}(V)\to B$ consisting of lines in the fibres of $V\to B$. The ordinary cohomology group $H^{*}(\mathbb{P}(V))$ of $\mathbb{P}(V)$ is generated by the relative hyperplane class $p:=c_{1}(\mathcal{O}_{\mathbb{P}(V)}(1))$ as an $H^{*}(B)$-algebra as follows:
@@
where $q$ is the Novikov variable associated with vertical lines. For $q\neq 0$, this ring decomposes into a direct sum of $r$ copies of $H^{*}(B)$. This decomposition must persist under the full quantum deformation, parametrized by the full Novikov variables of $\mathbb{P}(V)$ and the bulk-deformation parameter $\hat{\tau}\in H^{*}(\mathbb{P}(V))$. The question then arises as to whether each summand is isomorphic to the quantum cohomology of $B$. This paper aims to affirm this correspondence, i.e. $QH^{*}(\mathbb{P}(V))$ decomposes into a direct sum of $r$ copies of $QH^{*}(B)$. Furthermore, we show that the decomposition can be lifted to quantum cohomology $D$-modules.
@@ 1.1. Mirror theorem for non-split projective bundles
The $J$-function $J_{V}^{\lambda}$ is defined over the Novikov ring of $V$. In Theorem 1.1, we use the splitting of $\pi_{*}\colon H_{2}(\mathbb{P}(V),\mathbb{Z})\twoheadrightarrow H_{2}(B,\mathbb{Z})=H_{2}(V,\mathbb{Z})$ given by the class $p=c_{1}(\mathcal{O}(1))$ to identify Novikov variables for $V$ with those for $\mathbb{P}(V)$, see Notation 3.1.
@@
The quantum Riemann-Roch theorem of Coates-Givental [7] determines the $S^{1}$-equivariant $J$-function $J_{V}^{\lambda}$ from the genus-zero Gromov-Witten invariants of the base $B$ and the Chern classes $c_{i}(V)$. Therefore Theorem 1.1 determines all genus-zero Gromov-Witten invariants of $\mathbb{P}(V)$ from those of the base $B$. This recovers the genus-zero part of the results of Fan-Lee [16] and Fan [15], who showed that all genus [^1] The method in this paper is not applicable to higher-genus reconstruction, as it relies on the quantum Lefschetz principle [34] for convex bundles. Gromov-Witten invariants of $\mathbb{P}(V)$ can be reconstructed from those of the base $B$ and the Chern classes $c_{i}(V)$. On the other hand, in contrast to the split case, it is not easy to find a closed formula for $J_{V}^{\lambda}$ in general. We give a non-split example, the cotangent bundle of $\mathbb{P}^{n}$, in Example 3.6.
@@
Several techniques exist for computing the Gromov-Witten invariants of non-split projective bundles. Strangeway [52] provided a reconstruction algorithm for certain projective bundles over $\mathbb{P}^{n}$. The quantum splitting principle of Lee-Lin-Qu-Wang [39] proposes a reduction to the split case by blowing up the base. Fan [14, Theorem 0.1, Question 4] explored a relationship between the $(V,e_{\lambda}^{-1})$-twisted Gromov-Witten invariants of $B$ and the $(\mathcal{O}(-1),e_{\lambda}^{-1})$-twisted (or untwisted) invariants of $\mathbb{P}(V)$, under the opposite assumption that $V$ is globally generated. Recently, Hinault-Yu-Zhang-Zhang [23] (see also earlier announcements [35, 36] by Katzarkov-Kontsevich-Pantev-Yu) proposed a reconstruction method based on the decomposition theorem (Theorem 1.7 below). The initial condition required for their reconstruction and the explicit algorithm based on Birkhoff factorization are provided in §5.8.
@@ 1.2. Decomposition of quantum $D$-modules
An important property of quantum cohomology is that it forms a $D$-module over the parameter space. Roughly speaking, the quantum $D$-module $\operatorname{QDM}(X)$ of a smooth projective variety $X$ is the module $H^{*}(X)[z][\![Q,\tau]\!]$ equipped with the flat covariant derivatives $\nabla_{\tau^{i}}$, $\nabla_{z\partial_{z}}$, $\nabla_{\xi Q\partial_{Q}}$ with respect to the parameters $\tau$, $z$, $Q$ (see §2.2). Here $\tau\in H^{*}(X)$ represents the bulk parameter and $Q$ denotes the Novikov variable. These covariant derivatives, called the quantum connection, are defined in terms of the quantum multiplication. The quantum $D$-module is further equipped with a $z$-sesquilinear pairing $P$, induced from the Poincaré pairing and compatible with the quantum connection. The $I$- or $J$-functions appearing in the previous section are cohomology-valued solutions of the quantum $D$-module.
@@
The Mirror Theorem 1.1 establishes that the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ is the Fourier transform of the $S^{1}$-equivariant quantum $D$-module $\operatorname{QDM}_{S^{1}}(V)$ (see Theorem 1.12 below). By combining this result with the quantum Riemann-Roch (QRR) theorem of Coates-Givental [7], which connects the Gromov-Witten invariants of $V$ and $B$, we prove that $\operatorname{QDM}(\mathbb{P}(V))$ decomposes into a direct sum of $r$ copies of $\operatorname{QDM}(B)$ after an appropriate localization.
To compare the quantum $D$-modules of $\mathbb{P}(V)$ and $B$, we embed the Novikov ring of $B$ into a localization of the Novikov ring of $\mathbb{P}(V)$. Let $(q,Q)$ denote the Novikov variable of $\mathbb{P}(V)$, where $q$ corresponds to the class of a vertical line and $Q$ is the Novikov variable of the base $B$, identified via the splitting in Notation 3.1. Introducing an additional copy $Q_{B}$ of the Novikov variable for $B$, we define the following embedding [^2] See Remark 5.2 for the geometric interpretation of this embedding. of Novikov rings:
(1.1) $\mathbb{C}[\![Q_{B}]\!]\hookrightarrow\mathbb{C}(\!(q^{-1/r^{\prime}})\!)[\![Q]\!]\qquad Q_{B}^{d}\mapsto q^{-c_{1}(V)\cdot d/r}Q^{d}$
where $r^{\prime}$ is $r$ or $2r$ depending on the parity of $r$ (see (5.1)). Let $\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}$ denote the base change of $\operatorname{QDM}(\mathbb{P}(V))$ to the localized base $\mathbb{C}(\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$, and let $\operatorname{QDM}(B)_{\rm ext,loc}$ denote the base change of $\operatorname{QDM}(B)$ via the embedding (1.1).
[…]
@@
Figure 1. Decomposition pictured over the global Kähler moduli space: $\operatorname{QDM}(\mathbb{P}(V))$ splits into a direct sum of $\operatorname{QDM}(B)$ near the infinity divisor $(q^{-1/r^{\prime}}=0)$.
Theorem 1.7 (Theorem 5.1).
@@
Let $V\to B$ be as in Theorem 1.1. We have a formal invertible map $\varsigma=\bigoplus_{j=0}^{r-1}\varsigma_{j}\colon H^{*}(\mathbb{P}(V))\to\bigoplus_{j=0}^{r-1}H^{*}(B)$ over $\mathbb{C}(\!(q^{-1/r})\!)[\![Q]\!]$ and a decomposition of the quantum $D$-modules over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$:
$\Phi\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{i=0}^{r-1}\varsigma_{i}^{*}\operatorname{QDM}(B)_{\rm ext,loc}.$
The pairs $(q,Q)$ and $(q^{-1/r^{\prime}},Q_{B})$ of Novikov variables, related by $Q_{B}=q^{-c_{1}(V)/r}Q$, may be viewed as two coordinate charts on the global Kähler moduli space illustrated in Figure 1. The isomorphism $\Phi$ glues the quantum $D$-modules $\operatorname{QDM}(\mathbb{P}(V))$ and $\operatorname{QDM}(B)_{\rm ext}^{\boxplus r}=\operatorname{QDM}(B)[\![q^{-1/r^{\prime}}]\!]^{\boxplus r}$ over the intersection of their domains of definition — the localized base $\operatorname{Spf}\mathbb{C}(\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$ — to form a global $D$-module.
Corollary 1.8 (see §5.7 for the proof).
[…]
@@
In the split case, a similar decomposition of the quantum $D$-module (more generally, for toric bundles) follows from Brown’s theorem [3]: this has been discussed by Coates-Givental-Tseng [8] and Koto [38]. These previous works rely on the equivariant localization with respect to the $(\mathbb{C}^{\times})^{r-1}$-action (which exists only in the split case) and our argument is substantively different from them.
Remark 1.11 (see [28, 35, 30]).
Theorem 1.7 provides a decomposition of quantum $D$-modules at the formal level. Given that quantum cohomology is expected to have convergent structure constants, we anticipate that the mirror map $\varsigma_{j}$ and the isomorphism $\Phi$ in Theorem 1.7 can be analytified with respect to the Novikov variables $q,Q$ and the bulk parameter $\hat{\tau}\in H^{*}(\mathbb{P}(V))$. However, we should not expect $\Phi$ to be analytic in $z$; rather, its matrix entries should belong to $\mathcal{O}^{\rm an}[\![z]\!]$, where $\mathcal{O}^{\rm an}$ denotes the sheaf of analytic functions in $q,Q,\hat{\tau}$. This formal (in $z$) decomposition should lift to an analytic decomposition over an angular sector in the $z$-plane. Furthermore, we expect this analytic decomposition to relate — at specific parameter values — to the semi-orthogonal decomposition [46] of the derived category of coherent sheaves:
$D^{b}(\mathbb{P}(V))=\langle D^{b}(B)_{0},\dots,D^{b}(B)_{r-1}\rangle,\qquad D^{b}(B)_{i}=D^{b}(B)\otimes\mathcal{O}(i)$
via the $\widehat{\Gamma}-$integral structure.
@@ 1.3. Equivariant quantum cohomology and Fourier transformation
$\lambda,\mathbb{S}]=z\mathbb{S}.$
@@
We can rephrase the mirror theorem (Theorems 1.1, 3.4) as the Fourier duality between the quantum $D$-modules of $V$ and $\mathbb{P}(V)$. In this paper, we treat $\operatorname{QDM}_{S^{1}}(V)$ as a $D$-module over the infinite-dimensional base $H^{*}_{S^{1}}(V)$ (see §2.3 and §4.1.2 for further details).
@@ Acknowledgements
H.I. thanks Fumihiko Sanda, Yuuji Tanaka and Constantin Teleman for very useful discussions on shift operators and equivariant cohomology. We thank anonymous referees for many valuable comments. H.I. is supported by JSPS grant 16H06335, 20K03582, 21H04994 and 23H01073. Y.K. is supported by JSPS grant 21J23023.
@@ 2.1. Gromov-Witten invariants and quantum cohomology
A reference on Gromov-Witten invariants is made to [43]. Let $X$ be a smooth projective variety over $\mathbb{C}$. For $n\geq 0$ and $d\in H_{2}(X,\mathbb{Z})$, let $X_{0,n,d}$ denote the moduli stack of genus-zero, $n$-points, degree $d$ stable maps to $X$. This is a proper Deligne-Mumford stack carrying a virtual fundamental class $X_{0,n,d}]_{\rm vir}\in H_{2D}(X_{0,n,d},\mathbb{Q})$ of degree $D=\dim_{\mathbb{C}}X+c_{1}(TX)\cdot d+n-3$. Let $\operatorname{ev}_{i}\colon X_{0,n,d}\to X$ be the evaluation map at the $i$th marked point. The (genus-zero, descendant) Gromov-Witten invariants are the integrals of the form:
@@
Let $\operatorname{Eff}(X)\subset H_{2}(X,\mathbb{Z})$ denote the semigroup of effective curve classes. For a module $K$, we write K[[Q]]K[\![Q]\!] for the space of formal power series $\sum_{d\in\operatorname{Eff}(X)}k_{d}Q^{d}$ with $k_{d}\in K$. The formal variable $Q$ is called the Novikov variable. The module K[[Q]]K[\![Q]\!] has a natural ring structure when $K$ is a ring. The quantum cohomology will be defined over the ring $\mathbb{C}[\![Q]\!]$, which is called the Novikov ring.
@@
for $\alpha,\beta,\gamma\in H^{*}(X)$. We call $\tau$ the bulk(-deformation) parameter. The product $\star_{\tau}$ introduces an associative and supercommutative ring structure on $H^{*}(X)[\![Q,\tau]\!]=H^{*}(X)[\![Q]\!][\![\tau^{0},\dots,\tau^{s}]\!]$ such that $1\in H^{0}(X)$ is the unity. This is called quantum cohomology. The variables $\tau^{i}$ associated with odd classes $\phi_{i}$ have odd parity, and therefore anti-commute with each other by supercommutativity; more precisely we have $\tau^{i}\tau^{j}=(-1)^{|i||j|}\tau^{j}\tau^{i}$, $\tau^{i}\phi_{j}=(-1)^{|i||j|}\phi_{j}\tau^{i}$ with $|i|=\deg\phi_{i}$. The Novikov variable $Q$ is even and commutes with other elements. We also follow the Koszul sign convention when expanding the correlator $\left\langle\alpha,\beta,\gamma,\tau,\dots,\tau\right\rangle_{0,n+3,d}^{X}$ in power series of $\tau^{i}$. The parameter space $H^{*}(X)$ should be viewed as a formal supermanifold over $\mathbb{C}[\![Q]\!]$; we refer to [43] for the details.
@@ 2.2. Quantum connection and fundamental solution
The quantum connection admits the fundamental solution $M_{X}(\tau)$ defined as follows (see [19, §1], [47, Proposition 2], [25, Proposition 2.4], [6, Proposition 3.1] for the properties below):
@@
where $\{\phi^{j}\}$ is the dual basis of $\{\phi_{i}\}$ such that $(\phi_{i},\phi^{j})_{X}=\delta_{i}^{j}$. We used the String Equation [47, §1.2] in the second line. The other columns $M_{X}(\tau)\phi_{i}$ of $M_{X}(\tau)$ are given by the derivatives $z\partial_{\tau^{i}}J_{X}(\tau)$ of the $J$-function due to the differential equation (2.7).